# Extragradient methods with complexity guarantees for hierarchical variational inequalities

### v1

@@ 2.2 Hemi-variational inequalities
One can rewrite (2.3) as a monotone inclusion $0\in({\mathsf{F}}+\partial g)(z^{\ast})$. If $g=\operatorname{\iota}_{\mathcal{C}}$ for a closed convex set $\mathcal{C}\subset\mathcal{Z}$, we obtain from the above the variational inequality $\operatorname{VI}({\mathsf{F}},\mathcal{C})$:
@@
Let $\mathcal{C}\subset\operatorname{dom}(g)$ be a given compact subset of $\operatorname{dom}(g)$. A popular convergence measures for $\operatorname{HVI}({\mathsf{F}},g)$ is the localized gap function
@@
where $H^{({\mathsf{F}},g)}:\mathcal{Z}\times\mathcal{Z}\to[-\infty,\infty]$ is defined as $H^{({\mathsf{F}},g)}(z,y)\triangleq\langle{\mathsf{F}}(y),z-y\rangle+g(z)-g(y)$. Historically, this localized version of a gap function can be traced back to [Minty62] (see also [kinderlehrer2000introduction, NesDual07]). Lemma 2.2 below summarizes its role for the numerical resolution of the problem $\operatorname{HVI}({\mathsf{F}},g)$ under continuity assumptions on $\mathsf{F}$}.
@@ 2.3 Sharpness and error-bound property
Our geometric framework is phrased in terms of weak-sharpness of the lower-level solution set $\mathcal{S}_{2}$. Originally formulated for optimization problems in [burke1993weak], weak-sharpness in the context of variational inequalities has been defined in [patriksson1993unified]. Subsequently, implications in terms of error bounds of primal and dual gap functions have been stated in [marcotte1998weak, liu2016characterization]. The following definition of weak sharpness is from [Huang:2018aa].
@@
Based on [Huang:2018aa], we can give the following characterization of weak sharpness in terms of an error bound involving the dual gap function of $\operatorname{HVI}({\mathsf{F}},g)$.
@@
Consider problem $\operatorname{HVI}({\mathsf{F}},g)$ with $\operatorname{dom}(g)$ a closed convex nonempty subset of $\mathcal{Z}$. If $\mathcal{S}=\operatorname{zer}({\mathsf{F}}+\partial g)$ is weakly sharp, then
$(\forall z\in\mathcal{D}_{g}):\qquad\Theta(z|{\mathsf{F}},g,\operatorname{dom}(g))\geq\tau\operatorname{dist}(z,\mathcal{S}).$
Proof.
See Appendix B. $\blacksquare$
This Proposition shows that weak sharpness implies that the gap function satisfies an error bound property [pang1997error]. Motivated by this fact, we propose the following definition:
@@
Let $\mathcal{S}$ be the nonempty solution set of $\operatorname{HVI}({\mathsf{F}},g)$. We say $\mathcal{S}$ is $(\alpha,\rho)$-weak sharp with $\alpha>0$ and $\rho\geq 1$ if
$(\forall z^{\ast}\in\mathcal{S})(\forall z\in\mathcal{D}_{g}=\operatorname{dom}(g)):\qquad H^{({\mathsf{F}},g)}(z,z^{\ast})\geq\alpha\rho^{-1}\operatorname{dist}(z,\mathcal{S})^{\rho}.$ (2.5)
@@
An important class of examples arises when $\rho=1$, notably in monotone linear complementarity problems in finite dimensions under nondegeneracy conditions [burke1993weak, pang1997error]. In the case where the $\operatorname{HVI}({\mathsf{F}},g)$ reduces to a convex optimization problem, weak-sharpness implies an Hölderian errror bound, an assumption already imposed by [Cabot2005] in the context of hierarchical minimization. Specifically, let us assume that ${\mathsf{F}}=0$ and $g\in\Gamma_{0}(\mathcal{Z})$. Then $\mathcal{S}=\operatorname*{argmin}g$, and acordingly, $H^{(0,g)}(z,z^{*})=g(z)-\min g$ for $z^{*}\in\mathcal{S}$. Hence, (2.5) implies
@@
If $\rho=1$, this is the weak-sharpness condition of [burke1993weak]. The case $\rho=2$ corresponds to the "quadratic growth" condition of [drusvyatskiy2018error]. $\Diamond$
@@ 2.4 Constraint qualifications
We remark that Assumption 1 implies that $\mathcal{S}_{2}$ is a non-empty closed convex set. The next assumption is essentially a constraint qualification condition on the composite function
$x\mapsto f_{\bar{x}}(x)\triangleq\langle{\mathsf{F}}_{1}(\bar{x}),x-\bar{x}\rangle+g_{1}(x)-g_{1}(\bar{x})+\operatorname{\iota}_{\mathcal{S}_{2}}(x).$
@@
The solution set $\mathcal{S}_{1}$ of (P) satisfies
$\mathcal{S}_{1}=\operatorname{zer}({\mathsf{F}}_{1}+\partial g_{1}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}).$
Similar assumptions have been made in [Bot:2014aa] for solving constrained variational inequalities. Assumption 2 is motivated by conditions ensuring a non-smooth sum rule of
@@
Common assumptions ensuring this property are [BauCom16, Corollary 16.48]:
  • $\operatorname{int}\operatorname{dom}(g_{1})\cap\mathcal{S}_{2}\neq\varnothing$;
  • $\operatorname{dom}(g_{1})\cap\operatorname{int}(\mathcal{S}_{2})\neq\varnothing$;
  • if $\mathcal{Z}$ is finite-dimensional and $\operatorname{ri}\operatorname{dom}(g_{1})\cap\operatorname{ri}\operatorname{dom}(g_{2})\neq\varnothing$.
Since $\mathcal{S}_{2}\subseteq\operatorname{dom}(\partial g_{2})\subset\operatorname{dom}(g_{2})$, these assumptions ensure the computationally pleasant identity $\partial(g_{1}+g_{2})=\partial g_{1}+\partial g_{2}$.
2.5 Gap functions for hierarchical variational inequalities
Since we are searching for a solution of a hierarchical system of variational inequalities, we have to introduce different merit functions for measuring the feasibility and the optimality of a test point $z\in\mathcal{Z}$.
@@ 2.5 Gap functions for hierarchical variational inequalities
We note that if $z\in\mathcal{U}\cap\mathcal{S}_{2}$, and the regularity conditions stated in Lemma 2.2 hold, then by the same Lemma, the inequality $\Theta_{\rm Opt}(z|\mathcal{U}\cap\mathcal{S}_{2})\leq 0$ is equivalent to $z\in\mathcal{S}_{1}$. However, in general an inequality of the form $\Theta_{\rm Opt}(z|\mathcal{U}\cap\mathcal{S}_{2})\leq 0$, does not tell us much about the qualitative properties of the test point if $z\notin\mathcal{S}_{2}$. It is thus important to obtain a lower bound on the optimality gap.
@@
Suppose $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp. Then for all nonempty and compact subsets $\mathcal{U}_{2}\subseteq\operatorname{dom}(g_{2})$ and all $z\in\operatorname{dom}(g_{2})$, we have
@@
Proof.
See Appendix B. $\blacksquare$
Remark 2.2.
Lemma 2.8 allows us to establish a lower bound on the optimality gap over the solution set $\mathcal{S}_{2}$. It extends the corresponding result in [boct2025accelerating] to the operator case. [samadi_improved_2025] and [alves2025inertial] prove this result for special case of bilevel VIs on compact domains. $\Diamond$
@@ 3 Optimistic extragradient method
Our algorithmic design follows the popular Tikhonov regularization approach and employs a family of regularized problems with the data
${\mathsf{V}}_{\sigma}(z)\triangleq{\mathsf{F}}_{2}(z)+\sigma{\mathsf{F}}_{1}(z):\mathcal{Z}\to\mathcal{Z},\text{ and }G_{\sigma}(z)\triangleq g_{2}(z)+\sigma g_{1}(z)\in\Gamma_{0}(\mathcal{Z}).$
The parameter $\sigma\geq 0$ determines the relative importance of the upper level problem relative to the lower level problem. A decreasing sequence of regularization parameters $(\sigma_{k})_{k\geq 1}$ induces a corresponding sequence of operators ${\mathsf{V}}_{k}\triangleq{\mathsf{V}}_{\sigma_{k}}$ and convex functions $G_{k}\triangleq G_{\sigma_{k}}$. Being a sum of monotone and Lipschitz continuous mappings, ${\mathsf{V}}_{\sigma}$ is $L_{\sigma}\triangleq L_{{\mathsf{F}}_{2}}+\sigma L_{{\mathsf{F}}_{1}}$-Lipschitz and monotone. In the same vein, the combined mapping $G_{\sigma}$ is proper convex and lower semi-continuous. Furthermore, Assumption 1 guarantees the non-smooth sum rule $\partial G_{k}=\sigma_{k}\partial g_{1}+\partial g_{2}$. The conceptual algorithm we propose for solving (P) is described in Algorithm 1. The reader should consider this as a description of a method; Once concrete definitions for the step-size $(t_{k})_{k}$ and the regularization parameters $(\sigma_{k})_{k}$ are given, we obtain a bona-fide algorithm from the pseudo-code. Importantly, it should be observed that our algorithm requires only one evaluation of operators ${\mathsf{F}}_{1}$, ${\mathsf{F}}_{2}$ per iteration, in contrast to previous works [samadi_improved_2025, alves2025inertial].
Algorithm 1 Optimistic extragradient for hierarchical HVI’s (P)
$z^{1}=z^{1/2}$, step-size sequence $(t_{k})_{k\geq 1}$, non-increasing regularization sequence $(\sigma_{k})_{k\geq 1}$ satisfying
@@ 3.1 Statement of the main results
The main result of this paper are two complexity statements for the averaged iterates generated by Algorithm 1 in terms of the introduced gap functions. We work in a specific geometric setting, involving key quantities associated with the lower level problem $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$. Associated to the bifunction $H^{({\mathsf{F}}_{2},g_{2})}$, Appendix A defines the mapping
@@
This function encodes dual properties of the variational inequality, since it holds that (cf. eq. (A.1))
@@
The following summability condition is essentially due to [Attouch:2010aa, AttCzarPey11]:
@@
$(\forall p^{*}\in\text{Range}(\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})):\;\sum_{k=1}^{\infty}t_{k}\left[\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\right]<\infty.$ (3.4)
To understand the meaning of Assumption 3, it is instructive to specialize our setting to the simple bilevel optimization case (cf. Section 1.1.1). In that case, the data of the lower level problem are identified with ${\mathsf{F}}_{2}=\nabla f_{2}$ and $\mathcal{S}_{2}=\operatorname*{argmin}_{z}(f_{2}+g_{2})(z)$, and thus
$H^{({\mathsf{F}}_{2},g_{2})}(x,y)$ $\leq f_{2}(x)-f_{2}(y)+g_{2}(x)-g_{2}(y)\triangleq\hat{f}_{2}(x)-\hat{f}_{2}(y)$
$=(\hat{f}_{2}-\min\hat{f}_{2})(x)-(\hat{f}_{2}-\min\hat{f}_{2})(y).$
@@
In [AttCzarPey11, Peypouquet:2012aa, boct2025accelerating] the following summability condition is imposed
@@
We see that this condition is stronger than our condition (3.4).
@@
Assumption 3 looks quite daunting to verify, but as already observed in [boct2025accelerating], it fits very natural to the geometric setting of this paper. Indeed, let us assume that $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp, with $\alpha>0$ and $\rho>1$. According to Definition 2.5, for all $z\in\operatorname{dom}(g_{2})$ and for all $z^{*}\in\mathcal{S}_{2}$, we have $H^{({\mathsf{F}}_{2},g_{2})}(z,z^{*})\geq\alpha\rho^{-1}\operatorname{dist}(z,\mathcal{S}_{2})^{\rho}$. Hence, eq. (3.2) yields for all $z\in\mathcal{S}_{2}$
@@
Assumption 3 will allow us to prove that the sequence generated by the algorithm is bounded, an important step towards deriving convergence rates in terms of the averaged trajectory
@@
Consider problem (P). Let Assumptions 1-2 hold. Let additionally either Assumption 3 hold or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be compact. Consider the step size $(t_{k})_{k\geq 1}$ with $8t_{k}^{2}L_{k}^{2}\leq 1$ for all $k\geq 1$, and let $(\sigma_{k})_{k\geq 1}$ be a non-increasing sequence satisfying (3.1).
@@
    Consider a nonempty compact set $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$. Then, there exists a constant $C_{\mathcal{U}_{2}}>0$ for which
@@
    In particular, under (3.1) all weak accumulation points $\tilde{z}$ of $(\bar{z}^{k})$ satisfy $\Theta_{\rm Feas}(\tilde{z}|\mathcal{U}_{2})\leq 0$.
@@
    If the lower level solution set $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp, then
    $B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2T_{K}(\alpha/\rho)}+\frac{C_{\mathcal{U}_{2}}\sum_{k=1}^{K}t_{k}\sigma_{k}}{(\alpha/\rho)T_{K}}\right]^{1/\rho}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{T_{K}\sigma_{K}}.$ (3.8)
    In particular, all weak accumulation points $\tilde{z}$ of $(\bar{z}^{k})$ satisfy $\Theta_{\rm Opt}(\tilde{z}|\mathcal{U}_{1}\cap\mathcal{S}_{2})=0$.

### v2

@@ 2.2 Hemi-variational inequalities (HVI’s)
One can rewrite (2.3) as a monotone inclusion $0\in({\mathsf{F}}+\partial g)(z^{\ast})$, or equivalently the solution set is $\mathcal{S}=\operatorname{zer}({\mathsf{F}}+\partial g)$. If $g=\operatorname{\iota}_{\mathcal{C}}$ for a closed convex set $\mathcal{C}\subset\mathcal{Z}$, we recover the variational inequality problem $\operatorname{VI}({\mathsf{F}},\mathcal{C})$:
@@
Let $\mathcal{C}\subset\operatorname{dom}(g)$ be a given compact subset. A popular merit function for $\operatorname{HVI}({\mathsf{F}},g)$ is the localized (dual) gap function
@@
where $H^{({\mathsf{F}},g)}:\mathcal{Z}\times\mathcal{Z}\to[-\infty,\infty]$ is defined as
$H^{({\mathsf{F}},g)}(z,y)\triangleq\langle{\mathsf{F}}(y),z-y\rangle+g(z)-g(y).$ (2.5)
Historically, this localized version of a gap function can be traced back to [49, 36]. Lemma 2.1 below, whose proof is given in Appendix C, extends a result proved in [50], and summarizes its role for the numerical resolution of the problem $\operatorname{HVI}({\mathsf{F}},g)$ under continuity assumptions on $\mathsf{F}$}.
@@ 2.3 Sharpness and error-bound property
Our geometric framework is phrased in terms of weak sharpness of the lower-level solution set $\mathcal{S}_{2}$. Originally formulated for optimization problems in [15], weak sharpness in the context of variational inequalities has been defined in [54]. Subsequently, implications in terms of error bounds of primal and dual gap functions have been developed in [44, 40]. The following definition of weak sharpness is from [31].
@@
Based on [31], we give the following characterization of weak sharpness in terms of an error bound for the dual gap function of $\operatorname{HVI}({\mathsf{F}},g)$; see proof in Appendix C.
@@
Consider problem $\operatorname{HVI}({\mathsf{F}},g)$ with $\operatorname{dom}(g)$ a closed, convex, and nonempty subset of $\mathcal{Z}$. If the solution set $\mathcal{S}$ of $\operatorname{HVI}({\mathsf{F}},g)$ is weakly sharp, then
$(\forall z\in\operatorname{dom}(g)):\qquad\Theta(z|{\mathsf{F}},g,\operatorname{dom}(g))\geq\tau\operatorname{dist}(z,\mathcal{S}).$
Proposition 2.3 shows that weak sharpness implies that the gap function satisfies an error bound property [53], motivating the following definition:
@@
Let $\mathcal{S}$ be the solution set of $\operatorname{HVI}({\mathsf{F}},g)$. We say $\mathcal{S}$ is $(\alpha,\rho)$-weakly sharp with $\alpha>0$ and $\rho\geq 1$ if
$(\forall z^{\ast}\in\mathcal{S})(\forall z\in\operatorname{dom}(g)):\qquad H^{({\mathsf{F}},g)}(z,z^{\ast})\geq\frac{\alpha}{\rho}\operatorname{dist}(z,\mathcal{S})^{\rho}.$ (2.6)
@@
An important class of examples arises when $\rho=1$, notably in monotone linear complementarity problems in finite dimensions under nondegeneracy conditions [15, 53]. In the case where the $\operatorname{HVI}({\mathsf{F}},g)$ reduces to a convex optimization problem, weak sharpness implies an Hölderian error bound, an assumption already imposed by [1] in the context of hierarchical minimization. Specifically, let us assume that ${\mathsf{F}}=0$ and $g\in\Gamma_{0}(\mathcal{Z})$. Then $\mathcal{S}=\operatorname*{argmin}g$, and accordingly, $H^{(0,g)}(z,z^{*})=g(z)-\min g$ for $z^{*}\in\mathcal{S}$. Hence, (2.6) implies
@@
If $\rho=1$, this is the weak-sharpness condition of [15]. The case $\rho=2$ corresponds to the "quadratic growth" condition of [22], a relaxation of strong convexity. $\Diamond$
@@ 2.4 Constraint qualifications
We remark that Assumption 1 implies that $\mathcal{S}_{2}$ is a non-empty, closed, and convex set. The next assumption is essentially a constraint qualification condition on the composite function $x\mapsto\langle{\mathsf{F}}_{1}(\bar{x}),x-\bar{x}\rangle+g_{1}(x)-g_{1}(\bar{x})+\operatorname{\iota}_{\mathcal{S}_{2}}(x).$
@@
The solution set $\mathcal{S}_{1}$ of (P) satisfies $\mathcal{S}_{1}=\operatorname{zer}({\mathsf{F}}_{1}+\partial g_{1}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})\neq\varnothing$.
Similar assumptions have been made in [14] for solving constrained VI’s. Assumption 2 is motivated by conditions ensuring a non-smooth sum rule of the form
@@
Common assumptions ensuring this property can be found in [9, Corollary 16.48].
2.5 Gap functions for hierarchical HVI’s
Since we are searching for a solution of a hierarchical system of VIs, we have to introduce different merit functions for measuring the feasibility and the optimality of a test point $z\in\mathcal{Z}$.
@@ 2.5 Gap functions for hierarchical HVI’s
We note that if $z\in\mathcal{U}\cap\mathcal{S}_{2}$, and the regularity conditions stated in Lemma 2.1 hold, then by the same lemma, the inequality $\Theta_{\rm Opt}(z|\mathcal{U}\cap\mathcal{S}_{2})\leq 0$ is equivalent to $z\in\mathcal{S}_{1}$. However, in general an inequality of the form $\Theta_{\rm Opt}(z|\mathcal{U}\cap\mathcal{S}_{2})\leq 0$, does not tell us much about the qualitative properties of the test point if $z\notin\mathcal{S}_{2}$. It is thus important to obtain a lower bound on the optimality gap. This is achieved by Lemma 2.7 below, with proof provided in Appendix C. It extends the corresponding result in [13] to the operator case. [57] and [2] prove this result for the special case of bilevel VIs on compact domains.
@@
Suppose $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp. Then for all nonempty and compact subsets $\mathcal{U}_{2}\subseteq\operatorname{dom}(g_{2})$ s.t. $\mathcal{U}_{2}\cap\mathcal{S}_{2}\neq\varnothing$, and all $z\in\operatorname{dom}(g_{2})$, we have
@@ 3 Optimistic extragradient method
Our algorithmic design follows the popular Tikhonov regularization approach and, at iteration $k\geq 1$, uses the operators ${\mathsf{V}}_{k}$ and $G_{k}$ defined in (1.1), which combine the upper- and lower-level data. A decreasing sequence of regularization parameters $\sigma_{k}\geq 0$ determines the relative importance of the upper-level problem relative to the lower-level problem. Being a sum of monotone and Lipschitz continuous mappings, ${\mathsf{V}}_{k}$ is monotone and $L_{k}$-Lipschitz with $L_{k}\triangleq L_{{\mathsf{F}}_{2}}+\sigma_{k}L_{{\mathsf{F}}_{1}}$. The conceptual algorithm we propose for solving (P) is described in Algorithm 1. Once concrete definitions for the step-sizes $(t_{k})_{k\geq 1}$ and the regularization parameters $(\sigma_{k})_{k\geq 1}$ are given, we obtain a bona-fide algorithm. Importantly, it should be observed that our method only requires one evaluation of operators ${\mathsf{F}}_{1}$, ${\mathsf{F}}_{2}$ per iteration, in contrast to previous works [57, 2].
Algorithm 1 Optimistic extragradient method for hierarchical HVI’s (P)
@@
$\begin{split}&\lim_{k\to\infty}\sigma_{k}=0,\qquad\lim_{T\to\infty}\sum_{k=1}^{T}t_{k}=\infty,\\ &\lim_{T\to\infty}\sigma_{T}\sum_{k=1}^{T}t_{k}=\infty\text{, and }\lim_{T\to\infty}\frac{\sum_{k=1}^{T}t_{k}\sigma_{k}}{\sum_{k=1}^{T}t_{k}}=0.\end{split}$ (3.1)
@@
Remark 3.1.
To simplify the presentation, we implicitly assume that the proximal mapping of $G_{k}$ can be evaluated efficiently. In many practically relevant situations this is indeed the case. Trivially, if $g_{1}$ or $g_{2}$ is zero, we obtain an easier implementation. Other examples include the cases where either $g_{1}$ is an $L_{1}$ norm and $g_{2}$ is an indicator function of a simple set, such as a box or an $L_{2}$-ball [9]. In general though, the evaluation of the proximal operator of the composite mapping $G_{k}=\sigma_{k}g_{1}+g_{2}$ is not an easy task. We can circumvent this problem, by adopting a lifted formulation [21]. To obtain such a formulation, let $x=[z_{1},z_{2}]\in\mathcal{Z}\otimes\mathcal{Z}\triangleq\mathcal{X}$ and consider the operators
$\mathsf{A}_{1}(x)=\begin{pmatrix}\partial g_{1}(z_{1})\\  0\end{pmatrix},\mathsf{A}_{2}(x)=\begin{pmatrix}0\\  \partial g_{2}(z_{2})\end{pmatrix}:\mathcal{X}\to 2^{\mathcal{X}},\text{ and }$
$\mathsf{B}_{1}(x)=\begin{pmatrix}{\mathsf{F}}_{1}(z_{1})\\ 0\end{pmatrix},\mathsf{B}_{2}(x)=\begin{pmatrix}z_{1}-z_{2}\\  {\mathsf{F}}_{2}(z_{2})-z_{1}+z_{2}\end{pmatrix}:\mathcal{X}\to\mathcal{X}.$
It is easy to see that $\tilde{\mathcal{S}}_{2}=\operatorname{zer}(\mathsf{A}_{2}+\mathsf{B}_{2})=\{(z_{1},z_{2})\in\mathcal{Z}\times\mathcal{Z}|z_{1}=z_{2},\text{ and }z_{2}\in\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})\}.$ This means we can obtain $\mathcal{S}_{2}$ as the projection of points in $\tilde{\mathcal{S}}_{2}$ onto its first coordinate. Moreover, the problem
$\text{Find }\bar{x}=[\bar{z}_{1},\bar{z}_{2}]\in\mathcal{X}\text{ s.t. }\langle{\mathsf{F}}_{1}(\bar{z}_{1}),z_{1}-\bar{z}_{1}\rangle+g_{1}(z_{1})-g_{1}(\bar{z}_{1})\geq 0\qquad\forall\bar{x}\in\tilde{\mathcal{S}}_{2}$
shares the same solution as the non-lifted formulation (P) after projection. We can thus approach the solution of the lifted problem by performing the parallel computations
$x^{k+1/2}=(\operatorname{Id}_{\mathcal{X}}+t_{k}\mathcal{A}_{k})^{-1}(x^{k}-t_{k}\mathcal{B}_{k}(x^{k-1/2})),$
$x^{k+1}=(\operatorname{Id}_{\mathcal{X}}+t_{k}\mathcal{A}_{k})^{-1}(x^{k}-t_{k}\mathcal{B}_{k}(x^{k+1/2})),$
where $\mathcal{A}_{k}\triangleq\sigma_{k}\mathsf{A}_{1}+\mathsf{A}_{2}$ and $\mathcal{B}_{k}\triangleq\sigma_{k}\mathsf{B}_{1}+\mathsf{B}_{2}$. Thanks to the direct product structure of the lifting, it is easy to see that the resolvents indeed involve parallel computations on the individual factors. As an illustration, the update of $x^{k+1/2}=[z_{1}^{k+1/2},z_{2}^{k+1/2}]$ reads explicitly as
$z_{1}^{k+1/2}$ $=\operatorname{\textsf{prox}}_{t_{k}\sigma_{k}g_{1}}(z_{1}^{k}-t_{k}(z_{1}^{k-1/2}-z_{2}^{k-1/2}+\sigma_{k}{\mathsf{F}}_{1}(z_{1}^{k-1/2}))),$
$z^{k+1/2}_{2}$ $=\operatorname{\textsf{prox}}_{t_{k}g_{2}}(z_{2}^{k}-t_{k}(z_{2}^{k-1/2}-z_{1}^{k-1/2}+{\mathsf{F}}_{2}(z_{2}^{k-1/2})).$
$\Diamond$
@@ 3.1 Statement of the main results
The main result of this paper are two complexity statements for the averaged iterates generated by Algorithm 1 in terms of the gap functions (2.7) and (2.8). We work in a specific geometric setting in which we assume some structure on the solution set of the lower level problem $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$. Associated to the bifunction $H^{({\mathsf{F}}_{2},g_{2})}$, Appendix B introduces the mapping
@@
This function encodes dual properties of $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$, since it holds that (cf. eq. (B.1))
@@
The following summability condition is essentially due to [6, 5]:
@@
$(\forall p\in\text{Range}(\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})):\;\sum_{k=1}^{\infty}t_{k}\left[\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p)-\operatorname{\mathtt{s}}(\sigma_{k}p|\mathcal{S}_{2})\right]<\infty.$ (3.4)
To understand the meaning of Assumption 3, it is instructive to specialize our setting to the simple convex bilevel optimization case (cf. Section 1.1.1). In that case, we have ${\mathsf{F}}_{2}=\nabla f_{2}$ and $\mathcal{S}_{2}=\operatorname*{argmin}_{z}(f_{2}+g_{2})(z)$, and thus
$H^{({\mathsf{F}}_{2},g_{2})}(y,z)$ $\leq f_{2}(y)-f_{2}(z)+g_{2}(y)-g_{2}(z)\triangleq\hat{f}_{2}(y)-\hat{f}_{2}(z)$
$=(\hat{f}_{2}-\min\hat{f}_{2})(y)-(\hat{f}_{2}-\min\hat{f}_{2})(z).$
@@
In [5, 55, 13] the following summability condition is imposed
@@
We see that this condition is more restrictive than our condition (3.4).
@@
Assumption 3 looks quite daunting to verify, but as already observed in [13], it fits very natural to the geometric setting of this paper. Indeed, let us assume that $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp, with $\alpha>0$ and $\rho>1$. According to Definition 2.4, for all $z\in\operatorname{dom}(g_{2})$ and for all $z^{*}\in\mathcal{S}_{2}$, we have $H^{({\mathsf{F}}_{2},g_{2})}(z,z^{*})\geq\alpha\rho^{-1}\operatorname{dist}(z,\mathcal{S}_{2})^{\rho}$. Hence, eq. (3.3) yields for all $z\in\mathcal{S}_{2}$
@@
Next, we assume the non-smooth part $g_{1}$ of $\operatorname{HVI}({\mathsf{F}}_{1},g_{1})$ to satisfy a sort of finite variation property. Specifically, we define the variation over the set $\mathcal{A}\times\mathcal{B}\subseteq\mathcal{Z}\times\mathcal{Z}$ as
$|.\operatorname{\mathsf{Var}}(g_{1}|\mathcal{A}\times\mathcal{B})\triangleq\sup_{(x,y)\in\mathcal{A}\times\mathcal{B}}\absolutevalue{g_{1}(x)-g_{1}(y)}$.
Assumption 4.
The set $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is closed, and for any compact sets $\mathcal{A},\mathcal{B}\subset\mathcal{Z}$ satisfying $\mathcal{A},\mathcal{B}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$, we have $\operatorname{\mathsf{Var}}(g_{1}|\mathcal{A}\times\mathcal{B})<\infty$.
Assumption 3 will allow us to prove that the sequences generated Algorithm 1 are bounded, an important step towards deriving convergence rates in terms of the averaged trajectory
@@
Consider problem (P). Let Assumptions 1, 2, and 4 hold. Assume additionally that either Assumptions 3 hold, or that $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is compact. Let $(\bar{z}^{k})_{k\geq 1}$ as in (3.5) be generated by Algorithm 1 with $16t_{k}^{2}L_{k}^{2}\leq 1$ for all $k\geq 1$.
@@
    Let $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be non empty and compact. Then, there exists a constant $C_{\mathcal{U}_{2}}>0$ for which
@@
    If the lower level solution set $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp and $\mathcal{U}_{2}$ in (i) is s.t. $\mathcal{U}_{2}\cap\mathcal{S}_{2}\neq\varnothing$, then for the same set $\mathcal{U}_{1}$ as in (ii), we have the lower bound
    $-B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2T_{K}(\alpha/\rho)}+\frac{C_{\mathcal{U}_{2}}\sum_{k=1}^{K}t_{k}\sigma_{k}}{(\alpha/\rho)T_{K}}\right]^{\frac{1}{\rho}}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2}).$ (3.8)